Trying to figure out how to find absolute extreme values of a function on an interval.



treme values of the function on the intera Find the absolute extreme values of the function 5) f(O) = sin(8 +3,0 ses **
Answer #1

f(0) = sin(0+) f(o) = COS (0+ ) fC)=o ie, Cos (0+)-0 = cos 0+1I ie8=0 Cos (0+1)= Cos (25-) 0+ 1/2 = 211-11/ o= 21-11 = flo)